OpenStudy (anonymous):

Find the slant asymptote of the rational function. F(x)=4000+20x+0.0001x^2/x

OpenStudy (jmark):

Given F(x)=(4000+20x+0.0001x^2)/x long division 4000+20x+0.0001x^2/x =0.0001x+20 hence the slant asymptote =0.0001x+20 Know more about Horizontal Asymptote Calculator at http://goo.gl/RMdjSE